# Discrete homotopy and homology theories for finite posets

### v1

@@ 4.1. Discrete cubical homology
The case $p=0$ is trivial. The dimension of every maximal chain of $M$ is not greater than $1$, then this says $C_{p}^{\rm Simpl}(M)=0$ for $p>1$, so ${H}_{p}^{\rm Simpl}(\mathcal{K}(M))=0$ for $p>1$. It is not difficult to see that every $p$-cube with $p>1$ is degenerate, which means $C_{p}^{\rm Cube}(M)=0$. Then ${H}_{p}^{\rm Cube}(\mathcal{K}(M))=0$ for $p>1$.
Consider $p=1$. A nondegenerate $1$-cube can be represented by a directed edge $(v_{1},v_{2})$ of the Hasse diagram of $M$, while a nondegenerate $1$-simplex is $\{v_{1},v_{2}\}$ where $v_{1}$ and $v_{2}$ are vertices of $M$. Hence $C_{1}^{\rm Cube}(M)$ has the same basis as $C_{1}^{\rm Simpl}(M)$. Because the differentials $\partial_{1}^{\rm Cube}$ and $\partial_{1}^{\rm Simpl}$ are indeed identical, $\psi_{*}$ is an isomorphism for $p=1$. ∎
When $M$ is a poset with $|\mathcal{K}(M)|$ being an $m$-manifold, Proposition 4.5 would generalize to higher dimensions. Before proceeding further we need a useful excision property for finite posets.
@@
Let $M$ be a finite poset and let $L$ be a subposet of $M$. Then $|\mathcal{K}(M-L)|$ is a strong deformation retract of $|\mathcal{K}(M)|-|{\mathcal{K}(L)}|$.
@@
Let $M$ be a poset with $|\mathcal{K}(M)|$ being a closed $m$-manifold. Then the homomorphism $\psi_{*}:{H}_{p}^{\rm Cube}(M)\rightarrow{H}_{p}^{\rm Simpl}(\mathcal{K}(M))$ is a surjection for $p\not=m$.
@@
We prove the theorem by induction on the number $n$ of generators of chain complex $\mathcal{C}_{*}^{\rm Cube}(M)$. If $n=1$, then $M$ consists of a single vertex $v$. In each dimension $p$, there is exactly one $p$-cube $g:Q_{p}\rightarrow v$ of $M$ and exactly one ordered $p$-simplex $\{v,\cdots,v\}$ of $\mathcal{K}(M)$. Furthermore, $\psi(g)=\{v,\cdots,v\}$. Then $\psi$ is an isomorphism already on the chain level. Hence $\psi_{*}$ is surjective in this case.
Now suppose the theorem holds for any finite poset whose associated cubical chain complex has fewer than $n$ generators. Let $M$ be a poset with ${C}_{*}^{\rm Cube}(M)$ having $n$ generators. Given $p\geq 0$. If $p>m$, then
${H}_{p}^{\rm Simpl}(\mathcal{K}(M))=0,$
where the equality follows from [6, Lemma 3.27]. Therefore $\psi_{*}$ is surjective for $p>m$.
The case $p=0$ is trivial. Assume that $10<p\leq m-$ 1. Let $g:Q_{p}\rightarrow M$ be a generator of ${C}_{p}^{\rm Cube}(M)$. If $g(Q_{p})=M$, then
$H_{p}^{\rm Simpl}(\mathcal{K}(M))=H_{p}^{\rm Simpl}(\mathcal{K}(g(Q_{p})))=0$
since $g(Q_{p})$ is contractible, so we are done. Now assume $g(Q_{p})\not=M$. Consider the following commutative diagram, where $i_{*},j_{*}$ are induced by inclusion maps.
[…]
Note that $|\mathcal{K}(M-g(Q_{p}))|$ is an $m$-manifold, being a strong deformation retract of $|\mathcal{K}(M)|-|{\mathcal{K}(g(Q_{p}))}|$ by Lemma 4.6. It follows from the induction hypothesis that the map $\psi_{*}$ on the bottom line of the diagram is surjective for $p\leq m-1$. We shall show that $j_{*}$ is a surjection for $p\leq m-1$, which means that the map $\psi_{*}$ on the top line has the same property and the proof is complete.
Because $g(Q_{p})$ is contractible, ${\mathcal{K}(g(Q_{p}))}$ is collapsible [1, Theorem 4.2.11]. Let us take $N$ to be a regular neighbourhood of $|{\mathcal{K}(g(Q_{p}))}|$ in $|\mathcal{K}(M)|$. It is known that $N$ is a PL $m$-ball [11, Corollary 3.27]. By excision, we have
${H}_{*}^{\rm Sing}(|\mathcal{K}(M)|,|\mathcal{K}(M)|-|{\mathcal{K}(g(Q_{p}))}|)={H}_{*}^{\rm Sing}({\rm int}N,{\rm int}N-|{\mathcal{K}(g(Q_{p}))}|).$
@@
${H}_{*}^{\rm Sing}({\rm int}N,{\rm int}N-|{\mathcal{K}(g(Q_{p}))}|)$ $={H}_{*}^{\rm Sing}((\mathbb{R}^{m},\mathbb{R}^{m}-h(|{\mathcal{K}(g(Q_{p}))}|)),$
${H}_{*}^{\rm Sing}({\rm int}N-|{\mathcal{K}(g(Q_{p}))}|)$ $={H}_{*}^{\rm Sing}(\mathbb{R}^{m}-h(|{\mathcal{K}(g(Q_{p}))}|)).$
@@
${H}_{*}^{\rm Sing}(\mathbb{R}^{m}-h(|{\mathcal{K}(g(Q_{p}))}|)={H}_{*}^{\rm Sing}(\mathbb{R}^{m}-B^{p}).$
It follows that ${H}_{i}^{\rm Sing}(|\mathcal{K}(M)|,|\mathcal{K}(M-g(Q_{p}))|)=\mathbb{Z}$ for $i=m$ and vanishes otherwise. Hence
$H_{p}^{\rm Sing}(|\mathcal{K}(M-g(Q_{p}))|)=H_{p}^{\rm Sing}(|\mathcal{K}(M)|-|{\mathcal{K}(g(Q_{p}))}|).$
The long exact sequence of homology groups for the pair $(|\mathcal{K}(M)|,|\mathcal{K}(M-g(Q_{p}))|)$ then implies $j_{*}$ has the desired property. ∎
@@
Let $p\leq m-1$. Then $\psi_{*}:{H}_{p}^{\rm Cube}(M-g(Q_{p}))\rightarrow{H}_{p}^{\rm Simpl}(\mathcal{K}(M))$ is a surjection for any $g\in L_{p}(M)$ with $|\mathcal{K}(g(Q_{p}))|\subseteq{\rm int}|\mathcal{K}(M)|$, where $M$ is a poset with $|\mathcal{K}(M)|$ being an $m$-manifold.
Theorem 4.7 might suggest conjecturing that $H_{*}^{\rm Cube}(M)$ and $H_{*}^{\rm Simpl}(\mathcal{K}(M))$ are the same for all finite posets $M$. Now we construct an example showing this is not always the case.
@@
Let $M$ be the finite poset described below.
${e}$${f}$${c\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${d\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${a\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${b\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
@@
$H_{1}^{\rm Simpl}(\mathcal{K}(M))=0.$
Represent $1$-cubes $f:Q_{1}\rightarrow M$ by $(x_{0},x_{1})$, where $x_{i}=f(i)$, and represent $2$-cubes $f:Q_{2}\rightarrow M$ by $(x_{00},x_{10},x_{01},x_{11})$, where $x_{ij}=f(i,j)$. We shall prove $H_{1}^{\rm Cube}(M)\not=0$ by finding a $1$-cycle $\alpha\in C^{\rm Cube}_{1}(M)$ that is not a boundary. Define
@@
If $\beta\in C_{2}^{\rm Cube}(M)$ satisfies $\partial_{2}^{\rm Cube}(\beta)\not=0$, then $\beta$ is a linear combination
@@
If $\alpha$ is a boundary, there is a $\beta\in C_{2}^{\rm Cube}(M)$ such that $\partial_{2}^{\rm Cube}(\beta)=\alpha.$ Then we have
@@
As we have seen, for a given poset $M$, some information about the homology groups of $M$ can be squeezed out of Corollary 4.8 by computing discrete cubical homology groups of a subposet $M$ with fewer vertices. In this subsection, going a step further, we generalize Corollary 4.8 in the context of a class of finite posets.
Let $M$ be a homogeneous poset of dimension $m$. Define $M^{=p}$ as the subposet of elements of degree equal to $p$,
$M^{=p}=\{m^{p}_{1},\cdots,m^{p}_{i_{p}}|~{}m^{p}_{i_{j}}\in M,{\rm deg}(m^{p}_{i_{j}})=p\}.$
Label the vertices of $M$ by
(4.3) $M=\bigcup_{\begin{subarray}{c}1\leq t\leq i_{p}\\ 0\leq p\leq m\end{subarray}}m^{p}_{t}.$
It is no difficult to see that if $M$ is a poset with $|\mathcal{K}(M)|$ being an $m$-manifold, then $M$ is homogeneous.
Let $A=\mathcal{K}(\{m^{0}_{t_{0}},\cdots,m^{m}_{t_{m}}\})$ be an $m$-simplex of $\mathcal{K}(M)$ and let $I^{m}=[0,1]^{m}$ be the $m$-cube in $\mathbb{R}^{m}$. The last coordinate of $I^{m}$ is singled out for special reference and written as $I^{m}=I^{m-1}\times I$. Regard $|A|$ as a polyhedron of $I^{m}$ by the canonical homeomorphism $f$ given by
$f(m_{t_{i}})=\begin{cases}(0,0\cdots,0),&~{}\text{if}~{}i=0,\\ e_{i},&~{}\text{if}~{}1\leq i\leq m-1,\\ (\frac{1}{m},\frac{1}{m}\cdots,\frac{1}{m}),&~{}\text{if}~{}i=m.\end{cases}$
@@
Let $a,~{}a^{\prime}\in|\mathcal{K}(M)|$. We say $a$ is in the shadow of $a^{\prime}$ in $|\mathcal{K}(M)|$ if there is an $m$-chain $A$ of $M$ such that $|\mathcal{K}(A)|$ contains both $a$ and $a^{\prime}$, and $a$ is directly below $a^{\prime}$ in $|\mathcal{K}(A)|$.
@@
Let $X$ be a subpolyhedron of $|\mathcal{K}(M)|$. The shadow ${\rm sh}(X)$ of $X$ is the closure of the set of points of $X$ that are directly below some other point of $X$.
Given polyhedra $X\subseteq Y\subseteq|\mathcal{K}(M)|$ such that $Y\searrow X$ is an elementary collapse, we call this collapse sunny if no points of $Y-X$ are in ${\rm sh}(Y)$. If $X$ is a point, then $Y$ is called sunny collapsible.
@@
Let $M$ be a poset with $|\mathcal{K}(M)|$ being an $m$-manifold and $f\in L_{n}(M)$. If ${\dim}f(Q_{n})<m$, then $|\mathcal{K}(f(Q_{n}))|$ is suuny collapsible.
@@
Denote by $v$ the image of $(0,\cdots,0)$ under $\mathcal{K}(f)$. We shall construct inductively a decreasing sequence of subpolyhedra
$|\mathcal{K}(f(Q_{n}))|=X_{0}\supseteq X_{1}\supseteq\cdots\supseteq X_{k}=v,$
@@
    $X_{i}=|vK^{n-i-1}|$, where $K^{n-i-1}$ is a principle $(n-i-1)$-complex.
@@
    There is a sunny collapse $X_{i-1}\searrow X_{i}$.
Let $K^{n-1}=\{A\in\mathcal{K}(f(Q_{n}))~{}|~{}v~{}\text{is not a vertex of}~{}A\}$. Then $X_{0}=|vK^{n-1}|$.
Suppose we are given $,X_{i-1},K^{n-i}$, satisfying the two conditions, we have to construct $X_{i},~{}K^{n-i-1}$ and prove the two conditions. Define
@@
which is a principle complex since $K^{n-i}$ is principle. Let $X_{i}=|vK^{n-i-1}|$.
For each $(n-i)$-simplex $D\in K^{n-i}$, the closure ${\rm cl}(v|D|)$ is a ball and $|D|$ is a face. Collapsing each ${\rm cl}(v|D|)$ from ${\rm cl}(|D|)$, we obtain
$|vK^{n-i}|\searrow|vK^{n-i-1}|,$
which is a sunny collapse because no points of ${\rm sh}(X_{i-1})$ are removed. ∎
[12, Lemma 20] states that if $(B^{m},B^{q})$ is an $(m,q)$-ball pair of codimension $m-q\geq 3$, then $B^{m}\searrow B^{q}$. If $M$ is a contractible poset with $|\mathcal{K}(M)|$ being an $m$-manifold, then $|\mathcal{K}(M)|$ is a PL ball since a collapsible PL manifold is a PL ball [11, Corollary 3.27]. Note that $(|\mathcal{K}(M)|,|\mathcal{K}(f(Q_{n}))|)$ may be not a ball pair [12], since $\partial|\mathcal{K}(f(Q_{n}))|$ and ${\rm int}|\mathcal{K}(f(Q_{n}))|$ are not necessarily contained in $\partial|\mathcal{K}(M)|$ and ${\rm int}|\mathcal{K}(M)|$, respectively. Moreover, the codimension may be less than $3$. Thus we cannot apply [12, Lemma 20] to show that $|\mathcal{K}(M)|\searrow|\mathcal{K}(f(Q_{n}))|$ if ${\dim}f(Q_{n})<m$. However, the techniques in this proof can be used to obtain our result.
@@
Let $M$ be a contractible poset with $|\mathcal{K}(M)|$ being an $m$-manifold and $f\in L_{n}(M)$. Then $|\mathcal{K}(M)|\searrow|\mathcal{K}(f(Q_{n}))|$ if ${\dim}f(Q_{n})<m$.
@@
The sunny collapse in the proof of Lemma 4.12 is simplicial, written as
@@
Let $L_{i}$ be the polyhedron consisting of $|\mathcal{K}(f(Q_{n}))|$ and all points below $|K_{i}|$ in $|\mathcal{K}(M)|$. We will prove that
$|\mathcal{K}(M)|\searrow L_{0}\searrow L_{1}\searrow\cdots\searrow L_{r}\searrow|K|.$
Choose a triangulation $Q$ of $|\mathcal{K}(M)|$ containing a subdivision $Q_{0}$ of $L_{0}$. The collapsibility of $|\mathcal{K}(M)|$ allows us to collapse $Q$ to $Q_{0}$ simplicially from the top, in order of decreasing dimension of simplices, completing the first step. The last step is obvious. If ${\rm deg}f(0,\cdots,0)<m$, then $L_{r}=|K|$. If ${\rm deg}f(0,\cdots,0)=m$, then $L_{r}$ consists of $K$ joined by some single arcs formed by points below $f(0,\cdots,0)$ in $|\mathcal{K}(M)|$. Collapsing these arcs finishes this step.

### v2

@@ 4.1. Discrete cubical homology
The case $p=0$ is trivial. The dimension of every maximal chain of $X$ is not greater than $1$, then this says $C_{p}^{\rm Simpl}(X)=0$ for $p>1$, so ${H}_{p}^{\rm Simpl}(\mathcal{K}(X))=0$ for $p>1$. It is not difficult to see that every $p$-cube with $p>1$ is degenerate, which means $C_{p}^{\rm Cube}(X)=0$. Then ${H}_{p}^{\rm Cube}(\mathcal{K}(X))=0$ for $p>1$.
Consider $p=1$. A nondegenerate $1$-cube can be represented by a directed edge $(v_{1},v_{2})$ of the Hasse diagram of $X$, while a nondegenerate $1$-simplex is $\{v_{1},v_{2}\}$ where $v_{1}$ and $v_{2}$ are vertices of $X$. Hence $C_{1}^{\rm Cube}(X)$ has the same basis as $C_{1}^{\rm Simpl}(X)$. Because the differentials $\partial_{1}^{\rm Cube}$ and $\partial_{1}^{\rm Simpl}$ are indeed identical, $\psi_{*}$ is an isomorphism for $p=1$. ∎
When $X$ is a poset with $|\mathcal{K}(X)|$ being an $m$-manifold, Proposition 4.5 would generalize to higher dimensions. Before proceeding further we need a useful excision property for finite posets.
@@
Let $X$ be a finite poset and let $Y$ be a subposet of $X$. Then $|\mathcal{K}(X-Y)|$ is a strong deformation retract of $|\mathcal{K}(X)|-|{\mathcal{K}(Y)}|$.
@@
Let $X$ be a poset with $|\mathcal{K}(X)|$ being a closed $m$-manifold. Then the homomorphism $\psi_{*}:{H}_{p}^{\rm Cube}(X)\rightarrow{H}_{p}^{\rm Simpl}(\mathcal{K}(X))$ is a surjection for $p\not=m$.
@@
We prove the theorem by induction on the number $n$ of generators of chain complex $\mathcal{C}_{*}^{\rm Cube}(X)$. If $n=1$, then $X$ consists of a single vertex $v$. In each dimension $p$, there is exactly one $p$-cube $\sigma:Q_{p}\rightarrow v$ of $X$ and exactly one ordered $p$-simplex $\{v,\cdots,v\}$ of $\mathcal{K}(X)$. Furthermore, $\psi(\sigma)=\{v,\cdots,v\}$. Then $\psi$ is an isomorphism already on the chain level. Hence $\psi_{*}$ is surjective in this case.
Now suppose the theorem holds for any finite poset whose associated cubical chain complex has fewer than $n$ generators. Let $X$ be a poset with ${C}_{*}^{\rm Cube}(X)$ having $n$ generators. Let $p\geq 0$. If $p>m$, then
${H}_{p}^{\rm Simpl}(\mathcal{K}(X))=0,$
where the equality follows from [9, Lemma 3.27]. Therefore $\psi_{*}$ is surjective for $p>m$.
The case $p=0$ is trivial. Assume that $0<p\leq m-1$. Let $\sigma:Q_{p}\rightarrow X$ be a generator of ${C}_{p}^{\rm Cube}(X)$. If $\sigma(Q_{p})=X$, then
$H_{p}^{\rm Simpl}(\mathcal{K}(X))=H_{p}^{\rm Simpl}(\mathcal{K}(\sigma(Q_{p})))=0$
since $\sigma(Q_{p})$ is contractible, so we are done. Now assume $\sigma(Q_{p})\not=X$. Consider the following commutative diagram, where $i_{*},j_{*}$ are induced by inclusion maps.
${{H}_{p}^{\rm Cube}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi_{*}}$${{H}_{p}^{\rm Simpl}(|\mathcal{K}(X)|)}$${{H}^{\rm Cube}_{p}(X-\sigma(Q_{p}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i_{*}}$$\scriptstyle{\psi_{*}}$${{H}_{p}^{\rm Simpl}(|\mathcal{K}(X-\sigma(Q_{p}))|)\ignorespaces\ignorespaces\ignorespaces\ignorespaces.}$$\scriptstyle{j_{*}}$
Note that $|\mathcal{K}(X-\sigma(Q_{p}))|$ is an $m$-manifold, being a strong deformation retract of $|\mathcal{K}(X)|-|{\mathcal{K}(\sigma(Q_{p}))}|$ by Lemma 4.6. It follows from the induction hypothesis that the map $\psi_{*}$ on the bottom line of the diagram is surjective for $p\leq m-1$. We shall show that $j_{*}$ is a surjection for $p\leq m-1$, which means that the map $\psi_{*}$ on the top line has the same property and the proof is complete.
Because $\sigma(Q_{p})$ is contractible, $\mathcal{K}(\sigma(Q_{p}))$} is collapsible [1, Theorem 4.2.11]. Let us take $N$ to be a regular neighbourhood of $|{\mathcal{K}(\sigma(Q_{p}))}|$ in $|\mathcal{K}(X)|$. It is known that $N$ is a PL $m$-ball [16, Corollary 3.27]. By excision, we have
${H}_{*}^{\rm Sing}(|\mathcal{K}(X)|,|\mathcal{K}(X)|-|{\mathcal{K}(\sigma(Q_{p}))}|)={H}_{*}^{\rm Sing}({\rm int}N,{\rm int}N-|{\mathcal{K}(\sigma(Q_{p}))}|).$
@@
${H}_{*}^{\rm Sing}({\rm int}N,{\rm int}N-|{\mathcal{K}(\sigma(Q_{p}))}|)$ $={H}_{*}^{\rm Sing}((\mathbb{R}^{m},\mathbb{R}^{m}-h(|{\mathcal{K}(\sigma(Q_{p}))}|)),$
${H}_{*}^{\rm Sing}({\rm int}N-|{\mathcal{K}(\sigma(Q_{p}))}|)$ $={H}_{*}^{\rm Sing}(\mathbb{R}^{m}-h(|{\mathcal{K}(\sigma(Q_{p}))}|)).$
@@
${H}_{*}^{\rm Sing}(\mathbb{R}^{m}-h(|{\mathcal{K}(\sigma(Q_{p}))}|)={H}_{*}^{\rm Sing}(\mathbb{R}^{m}-B^{p}).$
It follows that ${H}_{i}^{\rm Sing}(|\mathcal{K}(X)|,|\mathcal{K}(X-\sigma(Q_{p}))|)=\mathbb{Z}$ for $i=m$ and vanishes otherwise. Hence
$H_{p}^{\rm Sing}(|\mathcal{K}(X-\sigma(Q_{p}))|)=H_{p}^{\rm Sing}(|\mathcal{K}(X)|-|{\mathcal{K}(\sigma(Q_{p}))}|).$
The long exact sequence of homology groups for the pair $(|\mathcal{K}(X)|,|\mathcal{K}(X-\sigma(Q_{p}))|)$ then implies $j_{*}$ has the desired property. ∎
@@
Let $p\leq m-1$. Then $\psi_{*}:{H}_{p}^{\rm Cube}(X-\sigma(Q_{p}))\rightarrow{H}_{p}^{\rm Simpl}(\mathcal{K}(X))$ is a surjection for any $\sigma\in L_{p}(X)$ with $|\mathcal{K}(\sigma(Q_{p}))|\subseteq{\rm int}|\mathcal{K}(X)|$, where $X$ is a poset with $|\mathcal{K}(X)|$ being an $m$-manifold.
The following examples will help us understand the content of this theorem.
@@
Let $X$ be the finite poset depicted below.
${c}$${b\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${a\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
Let us compute the $1$-st discrete cubical homology group $H_{1}^{\rm Cube}(X)$. Represent $1$-cubes $\sigma:Q_{1}\rightarrow X$ by $(x_{0},x_{1})$, where $x_{i}=f(i)$, and represent $2$-cubes $\sigma:Q_{2}\rightarrow X$ by $(x_{00},x_{10},x_{01},x_{11})$, where $x_{ij}=\sigma(i,j)$. It is easy to recognize $1$-cycles looking into this Hasse diagram. More precisely, every $1$-cycle must be a multiple of
(a,b)+(b,c)(a,c).(a,b)+(b,c)-(a,c).
By direct computation,
$\partial_{2}^{\rm Cube}(a,b,a,c)=(a,b)+(b,c)-(a,c),$
which follows that
$H_{1}^{\rm Cube}(X)=0.$
Theorem 4.7 then forces the $1$-st simplicial homology group of $\mathcal{K}(X)$ to vanish, namely,
$H_{1}^{\rm Simpl}(\mathcal{K}(X))=0.$
Theorem 4.7 might suggest conjecturing that $H_{*}^{\rm Cube}(X)$ and $H_{*}^{\rm Simpl}(\mathcal{K}(X))$ are the same for all finite posets $X$. Now we construct an example showing this is not always the case.
Example 4.10.
Let $X$ be the finite poset described below.
[…]
@@
$H_{1}^{\rm Simpl}(\mathcal{K}(X))=0.$
Represent $1$-cubes $\sigma:Q_{1}\rightarrow X$ by $(x_{0},x_{1})$, where $x_{i}=\sigma(i)$, and represent $2$-cubes $\sigma:Q_{2}\rightarrow X$ by $(x_{00},x_{10},x_{01},x_{11})$, where $x_{ij}=\sigma(i,j)$. We shall prove $H_{1}^{\rm Cube}(X)\not=0$ by finding a $1$-cycle $\alpha\in C^{\rm Cube}_{1}(X)$ that is not a boundary. Define
@@
If $\beta\in C_{2}^{\rm Cube}(X)$ satisfies $\partial_{2}^{\rm Cube}(\beta)\not=0$, then $\beta$ is a linear combination
@@
If $\alpha$ is a boundary, there is a $\beta\in C_{2}^{\rm Cube}(X)$ such that $\partial_{2}^{\rm Cube}(\beta)=\alpha.$ Then we have
@@
Remark 4.11.
It sometimes happens that computing discrete homology groups for a poset has the effect of simplifying the computation of the classical homology groups. This is exactly what Example 4.9 tells us. Furthermore, compared with the classical homology groups for a poset, the discrete ones sometimes have the advantage of collecting more imformation about this poset with nontrivial values.
@@
As we have seen, for a given poset $X$, some information about the homology groups of $X$ can be squeezed out of Corollary 4.8 by computing discrete cubical homology groups of a subposet $Y\subseteq X$ with fewer vertices. In this subsection, going a step further, we generalize Corollary 4.8 in the context of a class of finite posets.
Let $X$ be a homogeneous poset of dimension $m$. Define $X^{=p}$ as the subposet of elements of degree equal to $p$,
$X^{=p}=\{x^{p}_{1},\cdots,x^{p}_{i_{p}}|~x^{p}_{i_{j}}\in X,{\rm deg}(x^{p}_{i_{j}})=p\}.$
Label the vertices of $X$ by
(4.3) $X=\bigcup_{\begin{subarray}{c}1\leq t\leq i_{p}\\ 0\leq p\leq m\end{subarray}}x^{p}_{t}.$
It is no difficult to see that if $X$ is a poset with $|\mathcal{K}(X)|$ being an $m$-manifold, then $X$ is homogeneous.
Let $A=\mathcal{K}(\{x^{0}_{t_{0}},\cdots,x^{m}_{t_{m}}\})$ be an $m$-simplex of $\mathcal{K}(X)$ where $1\leq t_{j}\leq i_{j},$ and let $I^{m}=[0,1]^{m}$ be the $m$-cube in $\mathbb{R}^{m}$. The last coordinate of $I^{m}$ is singled out for special reference and written as $I^{m}=I^{m-1}\times I$. Regard $|A|$ as a subpolyhedron of $I^{m}$ by the canonical homeomorphism $f$ given by
$f(x^{j}_{t_{j}})=\begin{cases}(0,0\cdots,0),&~\text{if}~j=0,\\ e_{j},&~\text{if}~1\leq j\leq m-1,\\ (\frac{1}{m},\frac{1}{m}\cdots,\frac{1}{m}),&~\text{if}~j=m.\end{cases}$
@@
Let $a,~a^{\prime}\in|\mathcal{K}(X)|$. We say $a$ is in the shadow of $a^{\prime}$ in $|\mathcal{K}(X)|$ if there is an $m$-chain AA of $X$ such that $|\mathcal{K}(A)|$ contains both $a$ and $a^{\prime}$, and $a$ is directly below $a^{\prime}$ in $|\mathcal{K}(A)|$.
@@
Let $P$ be a subpolyhedron of $|\mathcal{K}(X)|$. The shadow ${\rm sh}(P)$ of $X$ is the closure of the set of points of $P$ that are directly below some other point of $P$.
[…]
To get a feeling for what a sunny collapse is about, let us look at the following example.
Example 4.14.
Let $X=\{x_{1}^{0},x_{2}^{0},x_{1}^{1},x_{2}^{1},x_{1}^{2}\}$ be a homogeneous poset of dimension $2$. Figure 2 shows its Hasse diagram and the geometric realization $|\mathcal{K}(X)|$. Consider the polyhedra $P_{i}\subseteq|\mathcal{K}(X)|$, $1\leq i\leq 3$, defined as follows:
$P_{1}=|\mathcal{K}(\{x_{1}^{0},x_{1}^{1},x_{2}^{1},x_{1}^{2}\})|,$
$P_{2}=|\mathcal{K}(\{x_{1}^{0},x_{1}^{1},x_{1}^{2}\})|,$
$P_{3}=|\mathcal{K}(\{x_{1}^{0},x_{1}^{1}\}\cup\{x_{1}^{1},x_{1}^{2}\})|.$
[…]
Figure 2. The poset $X$ and the geometric realization $|\mathcal{K}(X)|$.
[…]
Sunny collapses will turn out to be quite useful for proving our main theorem, though at first glance this property may seem just an idle technicality. The following lemma shows that there exists a class of posets which are sunny collapsible. Moreover, the fact that these posets are sunny collapsible becomes important later at key points in the proof of Lemma 4.16, which plays a fundamental role for obtaining our main theorem.
@@
Let $X$ be a poset with $|\mathcal{K}(X)|$ being an $m$-manifold and $\sigma\in L_{n}(X)$. If ${\dim}\sigma(Q_{n})<m$, then $|\mathcal{K}(\sigma(Q_{n}))|$ is sunny collapsible.
@@
Denote by $v$ the image of $(0,\cdots,0)$ under $\mathcal{K}(\sigma)$. We shall construct inductively a decreasing sequence of subpolyhedra
$|\mathcal{K}(\sigma(Q_{n}))|=P_{0}\supseteq P_{1}\supseteq\cdots\supseteq P_{k}=v,$
@@
    $P_{i}=|vK^{n-i-1}|$, where $K^{n-i-1}$ is a homogeneous (ni1)(n-i-1)-complex.
@@
[…]
Let $K^{n-1}=\{A\in\mathcal{K}(\sigma(Q_{n}))~|~v~\text{is not a vertex of}~A\}$. Then $P_{0}=|vK^{n-1}|$.
Suppose we are given $P_{i-1},~K^{n-i}$ satisfying the two conditions, we have to construct $P_{i},~K^{n-i-1}$ and verify the two conditions. Define
@@
which is a homogeneous complex since $K^{n-i}$ is homogeneous. Let $P_{i}=|vK^{n-i-1}|$.
For each (ni)(n-i)-simplex $D\in K^{n-i}$, the closure ${\rm cl}(v|D|)$ is a ball and $|D|$ is a face. Collapsing each ${\rm cl}(v|D|)$ from ${\rm cl}(|D|)$, we obtain
[…]
which is a sunny collapse because no points of ${\rm sh}(P_{i-1})$ are removed. ∎
[17, Lemma 20] states that if $(B^{m},B^{q})$ is an $(m,q)$-ball pair of codimension $m-q\geq 3$, then $B^{m}\searrow B^{q}$. If $X$ is a contractible poset with $|\mathcal{K}(X)|$ being an $m$-manifold, then $|\mathcal{K}(X)|$ is a PL ball since a collapsible PL manifold is a PL ball [16, Corollary 3.27]. Note that
$(|\mathcal{K}(X)|,|\mathcal{K}(\sigma(Q_{n}))|)$
may be not a ball pair [17], since $\partial|\mathcal{K}(\sigma(Q_{n}))|$ and ${\rm int}|\mathcal{K}(\sigma(Q_{n}))|$ are not necessarily contained in $\partial|\mathcal{K}(X)|$ and ${\rm int}|\mathcal{K}(X)|$, respectively. Moreover, the codimension may be less than $3$. Thus we cannot apply [17, Lemma 20] to show that
$|\mathcal{K}(X)|\searrow|\mathcal{K}(\sigma(Q_{n}))|$
if ${\dim}\sigma(Q_{n})<m$. However, the techniques thereof can be used to obtain this result.
@@
Let $X$ be a contractible poset with $|\mathcal{K}(X)|$ being an $m$-manifold and $\sigma\in L_{n}(X)$. Then $|\mathcal{K}(X)|\searrow|\mathcal{K}(\sigma(Q_{n}))|$ if ${\dim}\sigma(Q_{n})<m$.
@@
We write the sunny collapse in the proof of Lemma 4.15 as
@@
Let $W_{i}$ be the polyhedron consisting of $P_{0}$ and all points below $P_{i}$ in $|\mathcal{K}(X)|$. We will prove that
$|\mathcal{K}(X)|\searrow W_{0}\searrow W_{1}\searrow\cdots\searrow W_{r}\searrow P_{0}.$
Choose a triangulation $R$ of $|\mathcal{K}(X)|$ containing a subdivision $R_{0}$ of $W_{0}$. The collapsibility of $|\mathcal{K}(X)|$ allows us to collapse $R$ to $R_{0}$ simplicially from the top, in order of decreasing dimension of simplices, completing the first step. The last step is obvious. If ${\rm deg}\sigma(0,\cdots,0)<m$, then $W_{r}=P$. If ${\rm deg}\sigma(0,\cdots,0)=m$, then $W_{r}$ consists of $P$ joined by some single arcs formed by points below $\sigma(0,\cdots,0)$ in $|\mathcal{K}(X)|$. Collapsing these arcs finishes this step.